# Order Doesn’t Matter, But Reasoning Does: Training LLMs with Order-Centric Augmentation

### v1

@@ Abstract
Logical reasoning is essential for large language models (LLMs) to ensure accurate and coherent inference. However, LLMs struggle with reasoning order variations and fail to generalize across logically equivalent transformations. LLMs often rely on fixed sequential patterns rather than true logical understanding. To address this issue, we introduce an order-centric data augmentation framework based on commutativity in logical reasoning. We first randomly shuffle independent premises to introduce condition order augmentation. For reasoning steps, we construct a directed acyclic graph (DAG) to model dependencies between steps, which allows us to identify valid reorderings of steps while preserving logical correctness. By leveraging order-centric augmentations, models can develop a more flexible and generalized reasoning process. Finally, we conduct extensive experiments across multiple logical reasoning benchmarks, demonstrating that our method significantly enhances LLMs’ reasoning performance and adaptability to diverse logical structures. We release our codes and augmented data in https://anonymous.4open.science/r/Order-Centric-Data-Augmentation-822C/.
@@
Qianxi He1, Qianyu He1, Jiaqing Liang2†, Yanghua Xiao1† Weikang Zhou3, Zeye Sun3, Fei Yu3 1Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University 2School of Data Science, Fudan University 3Ant Group {qxhe23, qyhe21}@m.fudan.edu.cn, {liangjiaqing, shawyh}@fudan.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across various real-world applications Jaech et al. (2024); Dubey et al. (2024); Liu et al. (2024a). Logic reasoning Cummins et al. (1991) is essential for LLMs. It allows models to draw valid conclusions, maintain coherence, and make reliable decisions across tasks Pan et al. (2023); Liu et al. (2023).
@@
In fact, commutativity is a crucial property of logical reasoning. As established by Gödel’s completeness theorem Gödel (1930) and Tarski’s model theory Tarski (1956), commutativity means that independent logical units can be freely reordered without changing the essence of the logical structure. Therefore, in logical reasoning, first, independent premises are commutative. As shown in the upper half of Fig. 1, different orders of premises represent equivalent problem structures. Furthermore, as demonstrated by Gentzen’s proof theory Gentzen (1935), reasoning steps are also commutative, provided their dependencies are intact. As shown in the lower half of Fig. 1, changing the order of steps without disrupting the dependencies results in an equivalent reasoning process. However, altering the order of dependent steps disrupts inference and prevents a coherent path to the correct conclusion.
@@ 2.1 Order Effect of Language Models
Large language models are sensitive to reasoning order. While word order variations in natural language have little impact Cao et al. (2023); Abdou et al. (2022), disrupting the order in reasoning tasks significantly degrades performance. Chen et al. (2024) show that models perform optimally only when the premise order matches the sequence required for the reasoning process. To address this, Liu et al. (2024b) propose reorganizing premise order to reduce order sensitivity. However, this approach is task-specific and lacks generalizability. Furthermore, the Reversal Curse reveals that models fail to grasp logical equivalence when trained with a fixed linguistic order Berglund et al. (2023b). Golovneva et al. (2024) mitigate this by proposing reverse training, where LLMs learn both forward and reverse reasoning by randomly shuffling words or segments within a sentence. This highlights the need for diverse training data with varied orderings.
@@ 2.2 Logical Reasoning Enhancing
Symbolic reasoning enhances LLMs by transforming natural language into formal logic, providing a symbolic approach that helps models understand logic Olausson et al. (2023); Xu et al. (2024); Zhang et al. (2023). Training and inference strategies use adversarial pre-training, contrastive learning, and multi-step explicit planning to improve training efficiency and reasoning effectiveness Pi et al. (2022); Jiao et al. (2022); Zhao et al. (2023). ata augmentation creates diverse training and testing data, aiding models in generalizing better across logical tasks Han et al. (2022); Tafjord et al. (2021); Clark et al. (2020). LogiGLUE Luo et al. (2023) builds a large-scale logical benchmark through instruction fine-tuning across deductive, abductive, and inductive tasks. LogicBench Parmar et al. (2024) focuses on single-rule inference, evaluating LLMs on 25 reasoning patterns and exposing their weaknesses in complex reasoning and negation handling.
@@ 3 Problem Formulation
[…]
@@ 4.1.1 Shuffling the Order of Premises
For example, if the original order is $[P_{1},P_{2},P_{3},P_{4},\dots,P_{n}]$, after applying the permutation $\sigma$, the new order might be $[P_{3},P_{4},P_{1},P_{n},\dots,P_{2},\dots]$.
@@ 4.2.1 Leveraging LLMs for Logical Reasoning Solutions
Since logical reasoning datasets typically provide only a single label (e.g., true/false) without a Chain-of-Thought (CoT) reasoning process, we generate detailed step-by-step reasoning solutions to bridge this gap Xu et al. (2024). We use LLMs [^1] In our experiment, we use GPT-4o-mini. for this process. As shown in Fig. 3, the methodology consists of three main steps: (1) For datasets without First-Order Logic (FOL) expressions, We extract their premises and conclusion and convert them into the corresponding FOL representations. (2) The FOL-augmented premises, along with the ground truth labels, are fed into the model, prompting it to generate a step-by-step solution. Each step must clarify its purpose and reasoning, leading to a final conclusion. (3) The generated solutions are then reprocessed by the model to extract the premise indices and prerequisite step indices used in each reasoning step.
@@ 4.2.3 Generating Augmented Solution Sequences
Our goal is to generate a reasoning sequence that integrates all steps while maintaining dependency constraints, based on the principles outlined above. First, we identify all independent steps where $\mathcal{S}_{used}^{(i)}=\emptyset$ from the dataset, remove them from the DAG, and add them to the list of feasible step sequences List. Then, we iterate over all possible combinations of these steps to generate multiple different lists of valid sequences. Next, for each step still present in the DAG, we iterate through the steps of its $\mathcal{S}_{used}$. If $\mathcal{S}_{used}$ contains a step from the current List, we remove that step from $\mathcal{S}_{used}$.
@@ 5.2 Overall Performance
Tab. 5.1 shows that our method effectively improves model reasoning performance. Compared to Vanilla SFT, condition shuffling significantly enhances performance. LLaMA3-8B-Instruct achieves +11.05% and +11.5% improvements on RuleTaker’s sequential and shuffled evaluations, respectively. Similarly, Mistral-7B-Instruct-v0.3 shows improvements of 17.85% and 19.15% on the two LogicNLI test sets. Overall, the general performance gain ranges from 7% to 15%.
Similarly, adding answer step shuffling further improves performance over COT training. LLaMA3-8B-Instruct achieves +3.55% and +3.90% improvements on RuleTaker’s sequential and shuffled evaluations. Mistral-7B-Instruct-v0.3 shows improvements of +5.44% and +5.22% on FOLIO’s sequential and shuffled evaluations. Overall, the average performance gain is between 2% and 3%. It is worth noting that LLaMA3-8B-Instruct and LLaMA2-13B-Chat perform worse on the LogicNLI dataset in SFT with COT compared to Vanilla SFT. This could be due to the fact that COT does not always improve model performance across all tasks Liu et al. (2024c).
Additionally, LLaMA models generally perform 2-3% better on sequential evaluations than shuffled ones, highlighting their sensitivity to order. Nonetheless, condition and answer steps shuffling mitigates this effect. In contrast, Mistral-7B maintains stable performance across both settings, sometimes even outperforming sequential evaluations in shuffled tests.
@@ 6.1 Condition Augmentation with Varying Shuffling Degrees
Table 4: The performance of three different augmentation methods: the first row represents the original DAG-based Answer Steps Shuffled augmentation, the second row represents random step shuffling without dependencies in Sec. 6.2, and the third row represents the combined condition and answer augmentation method in Sec. 6.3.
@@ 6.2 The Importance of DAG-based Step Dependency
$TFI=\frac{\text{Number of valid sequences}}{\text{Factorial of the number of steps}}$ (1)
Number of valid sequences represents the number of valid topological orderings that respect the dependency constraints within the DAG. Factorial of the number of steps corresponds the number of possible orderings if no dependencies were present. The closer the TFI value is to 1, the looser the DAG structure, indicating higher reordering flexibility and greater step independence. Conversely, the closer the TFI value is to 0, the stronger the step dependencies, meaning the sequence must follow a strict order with little to no flexibility. Fig. 5 presents the TFI distribution across three datasets, illustrating the degree of step dependency in different reasoning tasks.
@@
In FOLIO, the majority of samples (49.9%) fall within the 0.0–0.1 range, indicating strong dependency constraints and minimal reordering flexibility. In contrast, RuleTaker and LogicNLI exhibit a significantly higher proportion of high-TFI samples (e.g., 44.0% and 27.0% in the 0.9–1.0 range, respectively), suggesting that these datasets contain more loosely connected reasoning structures. These trends highlight that FOLIO imposes stricter logical dependencies. In contrast, the latter datasets have weaker step dependencies, offering greater reordering flexibility. However, this may also imply that the quality of the generated COT needs improvement.
The difference in TFI across datasets aligns with the conclusions we obtained from the Random Step Shuffled experiment. Specifically, stronger step dependencies result in greater performance loss from random shuffling. Therefore, when performing answer order augmentation, it is important to maintain the integrity of these dependencies.
6.3 Combined Condition and Step Shuffling Leads to Performance Degradation
To investigate the combined effect of condition and step order augmentation, we apply an additional premise shuffle to Answer Steps Shuffled data, adjusting premise references in the answers accordingly. As shown in Tab. 4, Condition&Answer Shuffled results in lower performance compared to Answer Steps Shuffled alone.
We believe the key reason is that premise and step shuffling serve different learning purposes. Premise shuffling enables the model to recognize that independent conditions with commutativity can lead to the same answer, while step shuffling allows it to understand that different reasoning paths under the same condition can yield consistent conclusions. When applied separately, each augmentation enhances the model’s understanding of logical equivalence, thereby improving its overall reasoning ability. However, when condition and step shuffling are applied together, the logical structure is perturbed in two ways, requiring the model to simultaneously align different orders of both conditions and steps, increasing learning difficulty and reducing generalization. This suggests that excessive augmentation may introduce noise, making it harder for models to establish logical equivalence.
@@ 7 Conclusion
In this paper, we systematically study how to enhance the logical reasoning ability of LLMs by addressing their limitations in reasoning order variations. We introduce an order-centric data augmentation framework based on the principles of independency and commutativity in logical reasoning. Our method involves shuffling independent premises to introduce order variations and constructing directed acyclic graphs (DAGs) to identify valid step reorderings while preserving logical dependency. Extensive experiments across multiple logical reasoning benchmarks demonstrate that our method significantly improves LLMs’ reasoning performance and their adaptability to diverse logical structures.
@@ Appendix A Appendix
A.1 Details of Generating Solutions

### v2

@@ Abstract
Logical reasoning is essential for large language models (LLMs) to ensure accurate and coherent inference. However, LLMs struggle with reasoning order variations and fail to generalize across logically equivalent transformations. LLMs often rely on fixed sequential patterns rather than true logical understanding. To address this issue, we introduce an order-centric data augmentation framework based on commutativity in logical reasoning. We first randomly shuffle independent premises to introduce condition order augmentation. For reasoning steps, we construct a directed acyclic graph (DAG) to model dependencies between steps, which allows us to identify valid reorderings of steps while preserving logical correctness. By leveraging order-centric augmentations, models can develop a more flexible and generalized reasoning process. Finally, we conduct extensive experiments across multiple logical reasoning benchmarks, demonstrating that our method significantly enhances LLMs’ reasoning performance and adaptability to diverse logical structures. We release our codes and augmented data in https://github.com/qianxiHe147/Order-Centric-Data-Augmentation.
@@
Qianxi He1, Qianyu He1, Jiaqing Liang2† Weikang Zhou3, Zeye Sun3, Fei Yu3, Yanghua Xiao1† 1Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University 2School of Data Science, Fudan University 3Ant Group {qxhe23, qyhe21}@m.fudan.edu.cn, {liangjiaqing, shawyh}@fudan.edu.cn
footnotetext: Corresponding authors.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across various real-world applications Jaech et al. (2024); Dubey et al. (2024); Liu et al. (2024a). Logic reasoning Cummins et al. (1991) is essential for LLMs. It allows models to draw valid conclusions, maintain coherence, and make reliable decisions across tasks Pan et al. (2023); Liu et al. (2023a).
@@
In fact, commutativity is a fundamental property of logical reasoning, ensuring that independent logical units can be reordered without altering their meaning. As established by Gödel’s completeness theorem Gödel (1930) and Tarski’s model theory Tarski (1956), commutativity means that independent logical units can be freely reordered without changing the essence of the logical structure. Therefore, in logical reasoning, first, independent premises are commutative. As shown in the upper half of Fig. 1, different orders of premises represent equivalent problem structures. Furthermore, as demonstrated by Gentzen’s proof theory Gentzen (1935), reasoning steps are also commutative, provided their dependencies are intact. As shown in the lower half of Fig. 1, changing the order of steps without disrupting the dependencies results in an equivalent reasoning process. However, altering the order of dependent steps disrupts inference and prevents a coherent path to the correct conclusion.
@@ 2.1 Order Effect of Language Models
Large language models are sensitive to reasoning order (an example is shown in Appendix  A.1). While word order variations in natural language have little impact Cao et al. (2023); Abdou et al. (2022), disrupting the order in reasoning tasks significantly degrades performance. Chen et al. (2024) show that models perform optimally only when the premise order matches the sequence required for the reasoning process. To address this, Liu et al. (2024b) propose reorganizing premise order to reduce order sensitivity. However, this approach is task-specific and lacks generalizability. Furthermore, the Reversal Curse reveals that models fail to grasp logical equivalence when trained with a fixed linguistic order Berglund et al. (2023b). Golovneva et al. (2024) mitigate this by proposing reverse training, where LLMs learn both forward and reverse reasoning by randomly shuffling words or segments within a sentence. This highlights the need for diverse training data with varied orderings.
@@ 2.2 Logical Reasoning Enhancing
Symbolic reasoning enhances LLMs by transforming natural language into formal logic, providing a symbolic approach that helps models understand logic Olausson et al. (2023); Xu et al. (2024); Zhang et al. (2023). Training and inference strategies use adversarial pre-training, contrastive learning, and multi-step explicit planning to improve training efficiency and reasoning effectiveness Pi et al. (2022); Jiao et al. (2022); Zhao et al. (2023). Data augmentation creates diverse training and testing data, aiding models in generalizing better across logical tasks Han et al. (2022); Tafjord et al. (2021); Clark et al. (2020). LogiGLUE Luo et al. (2023) builds a large-scale logical benchmark through instruction fine-tuning across deductive, abductive, and inductive tasks. LogicBench Parmar et al. (2024) focuses on single-rule inference, evaluating LLMs on 25 reasoning patterns and exposing their weaknesses in complex reasoning and negation handling.
@@ 3 Problem Formulation
In this paper, we formulate the problem of logical reasoning in a unified representation. Let $D=\{P,C,L\}$ represent a logical reasoning problem, where $P=\{P_{1},P_{2},\dots,P_{n}\}$ is the set of premises, $C$ is the conclusion, and $L$ is the label, which takes a value from a finite set, such as $\{\text{true},\text{false},\text{uncertain}\}$, indicating whether $C$ can be logically inferred from $P$. In step-based data augmentation, we extend the representation to include a solution $S=\{S_{1},S_{2},\dots,S_{m}\}$, where $S$ consists of reasoning steps that derive the conclusion from the premises. This process can be abstracted as a directed acyclic graph (DAG). Typical logical reasoning datasets only provide labels. Therefore, we construct $S$ ourselves. The specific construction of $S$ will be detailed in Sec. 4.2.
@@ 4.1.1 Shuffling the Order of Premises
For example, if the original order is $P_{1},P_{2},P_{3},P_{4},\dots,P_{n}]$, after applying the permutation $\sigma$, the new order might be $P_{3},P_{4},P_{1},P_{n},\dots,P_{2},\dots]$.
@@ 4.2.1 Leveraging LLMs for Logical Reasoning Solutions
Since logical reasoning datasets typically provide only a single label (e.g., true/false) without a Chain-of-Thought (CoT) reasoning process, we generate detailed step-by-step reasoning solutions to bridge this gap Xu et al. (2024). We use LLMsIn our experiment, we use GPT-4o-mini. for this process. As shown in Fig. 3, the methodology consists of three main steps: (1) For datasets without First-Order Logic (FOL) expressions, We extract their premises and conclusion and convert them into the corresponding FOL representations. (2) The FOL-augmented premises, along with the ground truth labels, are fed into the model, prompting it to generate a step-by-step solution. Each step must clarify its purpose and reasoning, leading to a final conclusion. (3) The generated solutions are then reprocessed by the model to extract the premise indices and prerequisite step indices used in each reasoning step.
@@ 4.2.3 Generating Augmented Solution Sequences
Our goal is to generate a reasoning sequence that integrates all steps while maintaining dependency constraints, based on the principles outlined above. We represent the dependency graph as a dictionary, where each key is a step name and its value is a list of steps it depends on. First, we identify all independent steps where $\mathcal{S}_{used}^{(i)}=\emptyset$ from the dataset, remove them from the DAG, and add them to the list of feasible step sequences List. Then, we iterate over all possible combinations of these steps to generate multiple different lists of valid sequences. Next, for each step still present in the DAG, we iterate through the steps of its $\mathcal{S}_{used}$. If $\mathcal{S}_{used}$ contains a step from the current List, we remove that step from $\mathcal{S}_{used}$.
@@ 5.2 Overall Performance
Tab. 2 shows that our method effectively improves model reasoning performance. Compared to Vanilla SFT, condition shuffling significantly enhances results across multiple datasets, with overall gains ranging from 7% to 15%. Notably, it improves not only robustness to perturbed input order but also accuracy on original sequential test sets, suggesting enhanced logical reasoning ability rather than mere order tolerance. Further incorporating answer step shuffling brings additional improvements over CoT training, contributing an average boost of 2% to 3%, highlighting the benefit of diversified reasoning paths.
While LLaMA models tend to perform slightly better on sequential evaluations (typically 2–3% higher), our shuffling strategies help mitigate this gap.
@@ 5.2.1 Training Efficiency
Models Training Data Performance
GSM8K MATH
LLaMA3-8B-Instruct mathematics_trainset+FOLIO_trainset 65.28% 21.20%
+ FOLIO Condition Shuffled 65.81% 21.58%
+ FOLIO Answer Steps Shuffled 68.16% 22.56%
Table 3: Performance comparison of LLaMA3-8B-Instruct on GSM8K and MATH benchmarks when trained with different augmentation strategies.
5.2.2 Generalization
To validate whether our order-centric data augmentation on logical datasets can generalize to other reasoning domains, we conduct experiments on two mathematical datasets: GSM8K and MATH. Specifically, for these two datasets, we mix their training sets with the original FOLIO dataset, the condition-shuffled augmented dataset, and the answer-shuffled augmented dataset during training. The results in Tab. 3.
The results demonstrate that our method can effectively generalize to other reasoning tasks, improving their reasoning performance. We believe this provides promising evidence of the broader applicability of our approach.
We further analyze the impact of reasoning step counts and compare with prompt-based methods; see Appendix A.4 and A.5.
@@ 6.1 Condition Augmentation with Varying Shuffling Degrees
Condition&Answer Shuffled on Stronger LLM
Llama-3.3-70B-Instruct Answer Steps Shuffled 80.30% 78.08% 79.90% 80.00% 51.05% 49.65%
Condition&Answer Shuffled 78.82% (-1.48) 79.80% (+1.72) 83.95% (+4.05) 82.65% (+2.65) 52.00% (+0.95) 51.40% (+1.75)
Table 5: The performance of three different augmentation methods: the first row represents the original DAG-based Answer Steps Shuffled augmentation, the second row represents random step shuffling without dependencies in Sec. 6.2, and the third row represents the combined condition and answer augmentation method in Sec. 6.3. The last two rows show our new results using Llama-3.3-70B-Instruct model.
@@ 6.2 The Importance of DAG-based Step Dependency
[…]
The number of valid sequences reflects the count of topological orderings consistent with the DAG’s dependencies, while the factorial of step count represents all possible orderings without constraints. A TFI value near 1 indicates weak dependencies and high reordering flexibility; a value near 0 suggests strong dependencies and strict sequencing. Fig. 5 shows the TFI distribution across three datasets, highlighting variations in step dependency among reasoning tasks.
@@
In FOLIO, nearly half of the samples (49.9%) have low TFI (0.0–0.1), indicating strong step dependencies. In contrast, RuleTaker and LogicNLI have more high-TFI samples (44.0% and 27.0% in the 0.9–1.0 range), reflecting looser reasoning structures. These differences align with our Random Step Shuffling results—datasets with stronger dependencies suffer more from shuffling. Thus, preserving dependency integrity is crucial when applying answer order augmentation.
6.3 Combined Condition and Step Shuffling
To investigate the combined effect of condition and step order perturbations, we apply premise shuffling on top of the Answer Steps Shuffled data, adjusting premise references in the answers accordingly. As shown in Table 5, Condition&Answer Shuffled slightly underperforms compared to step-only shuffling in most cases.
This is likely because the two augmentations serve different purposes: condition shuffling teaches the model that independent premises are commutative, while step shuffling helps it understand that different reasoning paths under the same condition can lead to the same conclusion. When applied separately, each enhances logical equivalence learning; when combined, the dual perturbation increases learning difficulty, making it harder for smaller models to generalize.
To test this hypothesis, we evaluate the stronger LLaMA-3.3-70B-Instruct model. The results show a consistent 2–4% improvement from joint perturbation over step-only shuffling on most datasets, suggesting that larger models are better equipped to handle complex logical variations and capture underlying equivalence.
@@ 7 Conclusion
In this paper, we propose an order-centric data augmentation framework to improve LLMs’ logical reasoning by addressing their sensitivity to reasoning order. By leveraging principles of independency and commutativity, we shuffle independent premises and use DAGs to guide valid step reorderings. Experiments across multiple benchmarks show that our method enhances reasoning performance and adaptability to diverse logical structures.
@@ Appendix A Appendix
A.1 Premise Matter
From an overall performance perspective, in our main experiments, we observed that models generally perform worse on the shuffled testset compared to the sequential testset, as shown in the first row of results for each model in Tab. 2.
There is an examples in Tab. 7 that illustrate the importance of premise order. In the original order, the model correctly follows logical steps to reach the right conclusion. However, in the shuffled order, it fails to connect key premises, leading to an incorrect answer. This suggests that models may rely on surface-level cues rather than true logical understanding.
Original Premises: 1. Beethoven leads the Vienna Music Society. 2. Beethoven wrote Symphony No. 9. 3. Composers write music pieces. 4. Vienna Music Society premiered Symphony No. 9. 5. Symphony No. 9 is a music piece. 6. Orchestras are led by conductors. 7. Vienna Music Society is an orchestra. Conclusion: Some orchestras premiered music pieces. Label: True Model’s answer: Step 1: Identify the orchestra and the music piece. From premise 7, we know that Vienna Music Society is an orchestra: Orchestra(viennaMusicSociety) From premise 5, we know that Symphony No. 9 is a music piece: MusicPiece(symphony9) Step 2: Premiere of the music piece by the orchestra. From premise 4, we know that Vienna Music Society premiered Symphony No. 9: Premiered(viennaMusicSociety, symphony9) Step 3: Draw the conclusion. There exists an orchestra (Vienna Music Society) that premiered a music piece (Symphony No. 9). The conclusion is true. Shuffled Premises: 1. Symphony No. 9 is a music piece. 2. Composers write music pieces. 3. Beethoven wrote Symphony No. 9. 4. Vienna Music Society premiered Symphony No. 9. 5. Vienna Music Society is an orchestra. 6. Beethoven leads the Vienna Music Society. 7. Orchestras are led by conductors. Model’s answer: Step 1: Identify the orchestra and its leader. From premise 5, Vienna Music Society is an orchestra. From premise 6, Beethoven leads the Vienna Music Society. Step 2: Analyze the relationship between orchestras and music pieces Premise 3 states that composers write music pieces, but it does not mention orchestras premiering music pieces. The conclusion claims "Some orchestras premiered music pieces," but the premises do not provide information about which orchestras premiered music pieces. Step 3: Evaluate the conclusion Since the premises do not provide sufficient information to support the claim that "Some orchestras premiered music pieces," the conclusion cannot be confirmed.
Table 7: An example where the order of premises affects the reasoning outcome.
A.2 Details of Generating Solutions
@@
A.3 Kendall Tau Distance
The Kendall tau coefficient $\tau$ is calculated as follows:
$\tau=\frac{C-D}{\binom{n}{2}}$
For example, if the original premise order is $P_{1},P_{2},P_{3},\dots,P_{n}]$, a permutation function $\sigma$ might rearrange it to $P_{3},P_{1},P_{2},\dots,P_{n}]$. This process allows us to explore different levels of order perturbation, with the goal of analyzing how such variations affect model performance. Examples of premise orders corresponding to different $\tau$ values can be seen in Fig. 6.
Models Training LogicNLI Sequential LogicNLI Shuffled
LLaMA3-8B-Instruct SFT with COT 42.20% 41.65%
+ Answer Steps Shuffled 43.80% 42.80%
+ 3_Steps Shuffled 51.9% 50.15%
+ 4_Steps Shuffled 46.85% 44.95%
+ 5_Steps Shuffled 41.3% 40.85%
Mistral-7B-Instruct-v0.3 SFT with COT 46.40% 44.85%
+ Answer Steps Shuffled 47.35% 47.30%
+ 3_Steps Shuffled 52.25% 51.10%
+ 4_Steps Shuffled 45.90% 41.30%
+ 5_Steps Shuffled 41.85% 41.05%
Table 8: Performance on answer order augmentation LogicNLI subsets with different reasoning step lengths.
Models Method FOLIO
GPT-4 Logic-LM 78.92%
SymbCoT 82.33%
GPT-3.5-turbo Logic-LM 62.74%
SymbCoT 57.84%
COP 47.00%
LLaMA3-8B-Instruct + Condition Shuffled 70.44%
+ Answer Steps Shuffled 77.34%
Table 9: Performance comparison of different models and methods on the FOLIO benchmark. Our proposed Answer Steps Shuffled augmentation method achieves strong.
A.4 Impact of Reasoning Step Length in Answer Order Augmentation
To investigate the impact of reasoning step count on our method, we conduct experiments using the LogicNLI dataset, which features a relatively balanced distribution of reasoning chains with different step lengths. Specifically, there are 647 instances suitable for answer shuffling, from which we select 150 samples with step lengths of 3, 4, and 5. These samples are combined with 1,000 original CoT samples to perform answer-shuffling-based data augmentation. The results on LLaMA3-8B-Instruct and Mistral-7B-Instruct-v0.3 are shown in Tab. 8.
The results indicate that using only Step-3 answer shuffling unexpectedly outperforms the original Answer Steps Shuffled method. Additionally, as the number of steps increases from 3 to 5, the model’s performance gradually declines. This suggests that excessively complex reasoning step perturbations may negatively impact the model’s ability to learn logical commutativity and equivalence.
A.5 Comparison to Prompt_Based Methods
To further validate the effectiveness of our approach, we conduct a comparative study against three representative symbolic reasoning methods: Logic-LM Pan et al. (2023), SymbCoT Xu et al. (2024), and COP Liu et al. (2023b). These methods were primarily evaluated using proprietary large language models such as GPT-3.5-turbo and GPT-4. Among the datasets used in these works, FOLIO is included in our evaluation as well, enabling a direct performance comparison.
As shown in Tab.  9, our method, despite being trained on the smaller open-source model LLaMA3-8B-Instruct, consistently outperforms GPT-3.5-turbo across all three symbolic reasoning baselines. Although it does not surpass GPT-4, our approach achieves competitive results given the substantial gap in model scale and computational resources. These findings underscore the efficiency and effectiveness of our training framework, particularly in enhancing logical reasoning capabilities without reliance on proprietary or excessively large models.